# Impact of AI on Software Engineering Jobs

### v1

@@ Abstract.
Artificial Intelligence technology such as Large Language Models (LLMs) have become extremely popular in creating code. This has led to the conjecture that future software jobs will be exclusively conducted by LLMs, and the software industry will cease to exist. But software engineering is much more than producing code—notably, maintaining large software and keeping it reliable is a major part of software engineering, which LLMs are not yet capable of.
@@ 1. LLMs in Coding
While LLMs are suitable for pattern-based software engineering tasks—so essentially of a syntactic nature—, the capability of LLMs to capture and reason about program semantics remains unknown. This leads to a viewpoint that most noncritical software engineering will be automated in the future, with critical systems still involving developers to help reason about code. The opposite viewpoint could be to prevent LLMs from being used, at least for tasks involving program reasoning. Both of these viewpoints are extreme. The truth is somewhere in between—and could hold the key to future software job descriptions. This can allow us to understand the evolution of the software professional as we know them today.
@@ Acknowledgements.
The authors acknowledge discussions with attendees at the IFIP Working Group 2.4 meeting in Singapore held February 24-27, 2025. This work is partially supported by a Singapore Ministry of Education (MoE) Tier 3 grant MOE-MOET32021-0001. David Lo and Rohan Padhye provided useful pointers.

### v2

@@ Abstract.
Artificial Intelligence (AI) technology such as Large Language Models (LLMs) have become extremely popular in creating code. This has led to the conjecture that future software jobs will be exclusively conducted by LLMs, and the software industry will cease to exist. But software engineering is much more than producing code—notably, maintaining large software and keeping it reliable is a major part of software engineering, which LLMs are not yet capable of.
@@ 1. LLMs in Coding
While LLMs are suitable for pattern-based software engineering tasks, the capability of LLMs to capture and reason about program semantics remains unknown. This leads to a viewpoint that most noncritical software engineering will be automated in the future, with critical systems still involving developers to help reason about code. The opposite viewpoint could be to prevent LLMs from being used, at least for tasks involving program reasoning. Both of these viewpoints are extreme. The truth is somewhere in between—and could hold the key to future software job descriptions. This can allow us to understand the evolution of the software professional as we know them today.
@@ Acknowledgements.
The authors acknowledge discussions with attendees at the IFIP Working Group 2.4 meeting held February 24-27, 2025. This work is partially supported by a Singapore Ministry of Education (MoE) Tier 3 grant MOE-MOET32021-0001. David Lo and Rohan Padhye provided useful pointers.